# Don’t Overthink it. Preferring Shorter Thinking Chains for Improved LLM Reasoning

### v1

@@ 4.4 Ablation studies
We investigate two axis of short-m@k: the value of $m$ and the tie breaking method. For all experiments we use LN-Super-$49$B, reporting results over the three benchmarks described in Section 3.1. For the ablation studies we focus on controlling thinking compute.
We start by ablating different $m\in\{1,3,4,5,7,9\}$ for short-m@k. Results are shown in Figure 5(a). As observed in our main results, short-1@k outperforms others in low-compute regimes, while being less effective for larger compute budgets. Larger $m$ values seem to perform similarly, with higher $m$ values yielding slightly better results in high-compute scenarios.
Next, we analyze the tie-breaking rule of short-m@k. We suggest the selection of the shortest reasoning chain among the vote-leading options. We compare this strategy to random tie-breaking, and to tie breaking according to the longest reasoning chain among the options. As shown in Figure 5(b), the short-m@k strategy outperforms random tie-breaking. In contrast, choosing the option with the longest reasoning chain yields inferior results.
5 Finetuning using shorter trajectories
Drawing from our findings, we investigate whether fine-tuning on shorter reasoning chains enhances the accuracy of reasoning in LLMs. To do so, we follow the S1111 paradigm, which fine-tunes an LLM to perform reasoning using only $1,$$000$ trajectories (Muennighoff et al., 2025). We create three versions of the S1111 dataset, built from examples with the shortest, longest, and random reasoning chains among several generations. Below we describe the data generation and finetuning processes (Section 5.1), and then describe our results and main findings (Section 5.2).
5.1 Data creation and finetuning setup
To construct the three variants of S1111, we generate multiple responses for each S1111 question-answer pair. Specifically, for each example, we produce $10$ distinct answers using the QwQ-$32$B model, which we select for its superior performance among the evaluated models (Section 3). The generation process uses the hyperparameters detailed in Section 3.1. From these $10$ responses per example, we derive three dataset variants—S1111-short, S1111-long, and S1111-random—by selecting the shortest/longest/random response, respectively. This results in three datasets, each containing the same $1,$$000$ queries but with distinct reasoning trajectories and answers. A histogram of the reasoning token counts for each dataset variant is provided in Appendix B.
Following the S1111 approach, we finetune the Qwen-$2.5$-$32$B-Instruct model on the three S1111 variants. The finetuning hyperparameters are consistent with those used for the S1.11.11.11.1 model (Muennighoff et al., 2025), and training is conducted on $32$ H100100100100 GPUs. [^5] We match the number of gradient steps as in used for S1.1. Each model was finetuned for about 2 hours. The resulting models are evaluated using the benchmarks and experimental setup described in Section 3.1. Specifically, for each model we generate 20 answers per example, and report average accuracy.
5.2 Finetuning results
Results are presented in Table 3. For the AIME $2025$ and HMMT benchmarks, the S1111-short variant achieves superior performance while using fewer thinking tokens. While performance on AIME $2024$ is similar across models, S1111-short still demonstrates the shortest thinking. Aggregated results across benchmarks reveal that the S1111-short model improves relative performance by $2.8\%$ compared to the S1111-random baseline, with a reduction of $5.8\%$ in thinking tokens. Conversely, the S1111-long model consumes more tokens than S1111-random, but obtains similar performance.
These results suggest that training on shorter reasoning sequences can lead to models that not only exhibit reduced computational overhead but also enhanced performance. This observation aligns with our findings in Section 3, which shows that answers with shorter thinking trajectories tend to be more accurate. We believe that developing models that reason more effectively with less computation holds significant potential, as supported by related work (Lu et al., 2025).
Table 3: Results for our finetuned models over the S1111 variants: S1111-short/long/random. We report average accuracy over 20 generations per example. The S1111-short model improves performance over the other two models, while using fewer thinking tokens.
Metric AIME 2024 AIME 2025 HMMT Average
Thinking Tokens $\downarrow$ Acc. $\uparrow$ Thinking Tokens $\downarrow$ Acc. $\uparrow$ Thinking Tokens $\downarrow$ Acc. $\uparrow$ Thinking Tokens $\downarrow$ Acc. $\uparrow$
S1-random 16145 68.8 17798 59.3 19243 40.8 17729 56.3
S1-long 16912 67.3 17973 58.5 19397 42.1 18094 $(+2.1\%)$ 56.0
S1-short 15364 68.3 17195 60.2 17557 45.2 16706 $(-5.8\%)$ 57.9
6 Limitations and broader impact
The primary limitation of the short-m@k method is the reliance on batch decoding, as it requires parallel generation of multiple reasoning trajectories. This dependency might restrict its applicability in scenarios where inference memory is constrained. It should be noted that short-m@k can be used without batch decoding, although its efficiency gains will be lower. Additionally, while we show that finetuning on shorter reasoning chains can improve performance and efficiency, our experiments are limited to a specific model (Qwen-$2.5$-$32$B-Instruct) and dataset (S1).
In terms of broader impact, this work holds promise for enhancing the efficiency and accessibility of reasoning-LLMs by significantly lowering the required computational resources and time. By reducing these barriers, the technology could become more widely available, thereby democratizing access to advanced reasoning capabilities across a broader range of users and applications. However, as is often the case with advancements in efficiency, the decreased cost and increased scalability also carry the risk of enabling wider misuse or unintended applications of these powerful models.
7 Conclusion
In this work, we challenged the common assumption that increased test-time computation leads to better performance in reasoning LLMs. Through empirical analysis on three complex mathematical and reasoning benchmarks, we showed that shorter reasoning chains consistently outperform longer ones, both in accuracy and computational efficiency. Building on this insight, we introduced short-m@k, an inference method that prioritizes early-terminating generations. short-1@k, our most efficient variant, is preferred over traditional majority voting in low-compute settings. short-3@k, while slightly less efficient, outperforms majority voting across all compute budgets. To further validate our findings, we fine-tuned an LLM on short reasoning trajectories and observed improved accuracy and faster runtime, whereas training on longer chains yields diminishing returns. These findings highlight a promising direction for developing faster and more effective reasoning LLMs by embracing brevity over extended computation.
Acknowledgments and Disclosure of Funding
Appendix A Per benchmark results
We present the per-benchmark results for each of the criteria persented in Section 4.2. The sample-size ($k$) results are presented in Figures 6, 7 and 8. The thinking compute comparison results are presented in Figures 9, 10 and 11. The time-to-answer results per benchamrk are presented in Figures 12, 13 and 14.
@@ Appendix A Per benchmark results
Figure 6: AIME 2024 - sample size ($k$) comparison.
(a) LN-Super-$49$B
(b) R1111-$32$B
(c) QwQ-32B
Figure 7: AIME 2025 - sample size ($k$) comparison.
(a) LN-Super-$49$B
(b) R1111-$32$B
(c) QwQ-32B
Figure 8: HMMT Feb 2025 - sample size ($k$) comparison.
(a) LN-Super-$49$B
(b) R1111-$32$B
(c) QwQ-32B
Figure 9: AIME 2024 - thinking compute comparison.
(a) LN-Super-$49$B
(b) R1111-$32$B
(c) QwQ-32B
Figure 10: AIME 2025 - thinking compute comparison.
(a) LN-Super-$49$B
(b) R1111-$32$B
(c) QwQ-32B
Figure 11: HMMT Feb 2025 - thinking compute comparison.
(a) LN-Super-$49$B
(b) R1111-$32$B
(c) QwQ-32B
Figure 12: AIME 2024 - time-to-answer comparison.
(a) LN-Super-$49$B
(b) R1111-$32$B
(c) QwQ-32B
Figure 13: AIME 2025 - time-to-answer comparison.
(a) LN-Super-$49$B
(b) R1111-$32$B
(c) QwQ-32B
Figure 14: HMMT Feb 2025 - time-to-answer comparison.
Appendix B Thinking token histograms of S1 variants
Figure 15 shows the thinking token count histograms for the three variants of the S1 dataset (short/long/random) presented in Section 5.

### v2

@@ Appendix C Ablation studies
Appendix D Small models results
We present our main results (Sections˜3 and 4) using smaller models. We use Llama-$3.1$-Nemotron-Nano-88B-v11 [LN-Nano-88B; Bercovich and others, 2025] and R11-Distill-Qwen-77B [R11-77B; Guo et al., 2025]. Table˜7 (corresponds to Table˜1), presents a comparison between shortest/longest/random generation per example for the smaller models. As observed for larger models, using the shortest answer outperforms both random and longest answers across all benchmarks and models.
Table 7: Comparison between taking the shortest/longest/random generation per example.
GPQA-D AIME 2024 AIME 2025 HMMT Math Average
Thinking Tokens $\downarrow$ Acc. $\uparrow$ Thinking Tokens $\downarrow$ Acc. $\uparrow$ Thinking Tokens $\downarrow$ Acc. $\uparrow$ Thinking Tokens $\downarrow$ Acc. $\uparrow$ Thinking Tokens $\downarrow$ Acc. $\uparrow$
LN-Nano-8B
random 7003 52.2 10380 62.1 11869 46.5 12693 34.0 11647 47.5
longest 10594 $(+51\%)$ 41.4 16801 40.0 17140 33.3 18516 23.3 17486 $(+50\%)$ 32.2
[…]
R1-7B
random 7015 35.5 11538 57.8 12377 42.2 14693 25.0 12869 41.7
longest 11863 $(+69\%)$ 29.8 21997 26.7 21029 26.7 23899 13.3 22308 $(+73\%)$ 22.2
shortest 3438 $(-51\%)$ 46.5 05217 76.7 06409 53.3 06950 43.3 6192 $(-52\%)$ 57.8
Next, we analyze the performance of short-m@k using smaller models (see details in Section˜4). Figure˜19, Figure˜20 and Figure˜21 are the sample-size/compute/time-to-answer results for the small models over the math benchmarks, respectively. Figure˜22, Figure˜23 and Figure˜24 are the sample-size/compute/time-to-answer results for the small models for GPQA-D, respectively.
The performance of short-m@k using small models remain consistent with those observed in larger ones. short-1@k demonstrates a performance advantage over majority voting in low compute regimes, while short-3@k dominates it across all compute budgets.
(a) LN-Nano-8B
(b) R11-77B
Figure 19: Small models - sample size ($k$) comparison over math benchmarks.
(a) LN-Nano-8B
(b) R11-77B
Figure 20: Small models - thinking compute comparison over math benchmarks.
(a) LN-Nano-8B
(b) R11-77B
Figure 21: Small models - time-to-answer comparison over math benchmarks.
(a) LN-Nano-8B
(b) R11-77B
Figure 22: Small models - sample size ($k$) comparison over GPQA-D.
(a) LN-Nano-8B
(b) R11-77B
Figure 23: Small models - thinking compute comparison over GPQA-D.
(a) LN-Nano-8B
(b) R11-77B
Figure 24: Small models - time-to-answer comparison over GPQA-D.
Appendix E GPQA-D sequential results
The results for GPQA-D when accounting for sequential compute are presented in Figure˜25.
@@ Appendix E GPQA-D sequential results
(d) R1-670B
Figure 25: Comparing different methods for the GPQA-D benchmark under sequential compute.
Appendix F Backtracks under controlled length results
Below we present the results for the backtrack count under controlled length scenarios (Section˜5.2). The results over the math benchmarks are presented in Table˜8 and for GPQA-D in Table˜9.
Table 8: Average number of backtracks for correct (C), incorrect (IC) answers, binned by thinking length. Results are averaged across math benchmarks.
Model Thinking Tokens 0-5k 5-10k 10-15k 15-20k 20-25k 25-30k 30-32k
C/IC C/IC C/IC C/IC C/IC C/IC C/IC
LN-Super-49B 35/064 100/133 185/236 261/299 307/320 263/323 00/0304
R1-32B 29/074 088/166 171/279 244/351 334/370 268/355 326/1006
QwQ-32B 50/148 120/174 194/247 285/353 354/424 390/476 551/0469
R1-670B 58/027 100/086 143/184 222/203 264/254 309/289 352/0337
Table 9: Average number of backtracks for correct (C), incorrect (IC) answers, binned by thinking length. Results are reported for GPQA-D.
Model Thinking Tokens 0-5k 5-10k 10-15k 15-20k 20-25k 25-30k 30-32k
C/IC C/IC C/IC C/IC C/IC C/IC C/IC
LN-Super-49B 38/52 175/164 207/213 00/$0$ 00/$0$ 00/$0$ 00/$0$
R1-32B 39/54 194/221 301/375 525/668 00/$0$ 00/$0$ 00/$0$
QwQ-32B 65/71 169/178 333/358 378/544 357/703 00/$0$ 00/$0$
R1-670B 44/72 093/155 178/232 297/300 341/341 463/382 553/477
Appendix G Further details and results for finetuning experiments
The generation process of all three variants of S11 uses the hyperparameters detailed in Section˜3.1. Figure˜26 shows the thinking token count histograms for the three variants of the S1 dataset (short/long/random) presented in Section˜6.
As for finetuning, we follow the S11 approach, and finetune the Qwen-2.52.5-77B-Instruct and Qwen-2.52.5-3232B-Instruct models on the three S11 variants. The finetuning hyperparameters are consistent with those used for the S1.11.1 model (Muennighoff et al., 2025), and training is conducted on $32$ H100100 GPUs. [^7] We match the number of gradient steps as in used for S1.1. The resulting models are evaluated using the benchmarks and experimental setup described in Section˜3.1. Specifically, for each model we generate 20 answers per example, and report average accuracy.
Results for the 7B version model are presented in Table˜10.
@@ Appendix G Further details and results for finetuning experiments
Table 10: Results for our finetuned models over the S11 variants using Qwen-2.5-7B-Instruct: S11-short/long/random.
GPQA-D AIME 2024 AIME 2025 HMMT Math Average
Thinking Tokens $\downarrow$ Acc. $\uparrow$ Thinking Tokens $\downarrow$ Acc. $\uparrow$ Thinking Tokens $\downarrow$ Acc. $\uparrow$ Thinking Tokens $\downarrow$ Acc. $\uparrow$ Thinking Tokens $\downarrow$ Acc. $\uparrow$
S1-random 14095 39.1 25207 22.0 23822 22.0 25028 10.8 24686 18.2
[…]